# Operad of Posets 101: The Wixárika Posets

### v2

@@ 4. Algebras over the operad of Wixárika posets
A more detailed definition of operad can be found in  [STA04, CHA07, LV12, VAL14, NAO20, BRA21, HM22].
Example 29.
@@
Example 30.
Consider the tree grafting operad: Let $T(n)$ be the set of trees with $1$ root and $n$ leaves labeled (arbitrarily) $1$ through $n$. The collection $\{T(n)\}$ forms an operad by grafting the root of $g$ to the leaf of $f$ labeled $i$, as in Figure 4.
Figure 4. Operations on the operad tree (Image created by Wiki_cies, https://commons.wikimedia.org/w/index.php?title=User:Wiki_cies&action=edit&redlink=1).
Definition 31.
@@
Definition 32.
The Wixárica operad is the operad $W$ generated by points under the binary operation $\ast$ and the unary operation $D$.
Example 33.
@@
Figure 5. Left: The identity operation, the operation $D$ and the operation $\ast$. Right: A tree representing the element in Figure 3.
The composition of posets then allows us to see Wixárika posets as algebras over the operad $W$.
Example 34.
Consider Figure 6. On the left side we have at the top three copies of the posets $\langle 1\rangle$. Below is a sequence of transformations using $D$ and *. At the bottom we have the final result, a Wixárika poset. On the right side we have the corresponding tree of operations in which the leaves should be thought of as decorated by the poset $\langle 1\rangle$. The vertices with two incoming edges and one outgoing edge are labeled with * and the vertices with one incoming edge and one outgoing edge (the bird nest) are labeled with $D$.
Figure 6. To the left bottom we have a Wixárika poset, as an element of the algebra. To the right we have the tree of operations, an element of the operad of Wixárika posets. When this element is evaluated on six points it produces the Wixárika poset of the left. As the drops of water run down the tree, the poset on the left is being assembled.
Definition 35.
@@
Given $P$ a Wixárika poset, define
$D(\mathfrak{Z}\left(P\right)):=\mathfrak{Z}\left(D(P)\right).$
Proposition 36.
@@
Put together Definition 19 and Proposition 22 to get:
$\mathfrak{Z}\left(P\ast Q\right)=\mathfrak{Z}\left(P\right)(1-x)\mathfrak{Z}\left(Q\right)=\mathfrak{Z}\left(P\right)\ast\mathfrak{Z}\left(Q\right).$
By Proposition 23 and its Corollary 24 we have:
$\mathfrak{Z}\left(P\sqcup Q\right)=\mathfrak{Z}\left(P\right)\sqcup\mathfrak{Z}\left(Q\right).$
We obtain the explicitly formula
$D(\mathfrak{Z}\left(P\right))=\mathfrak{Z}\left(1\right)\ast(\mathfrak{Z}\left(P\right)\sqcup\mathfrak{Z}\left(1\right))\ast\mathfrak{Z}\left(1\right).$
5. Evaluation on order series
In this section, we give an explicit example of how to compute the order series of a Wixárica poset using the algebra homomorphims between both $W$-algebras defined in the previous section. To compute the order series, we follow [FS86], but work on the level of power series instead of order polynomials.
Consider the homomorphism between the algebra over the operad $W$ generated by $\langle 1\rangle$ and the algebra generated by $\mathfrak{Z}\left(1\right)$. In Proposition 36 we constructed $\ast,\sqcup$ such that
$\mathfrak{Z}\left(P\right)*\mathfrak{Z}\left(Q\right)=\mathfrak{Z}\left(P*Q\right),\quad D(\mathfrak{Z}\left(P\right))=\mathfrak{Z}\left(D(P)\right).$
@@ 5. Evaluation on order series
We shall give a procedure to calculate $\mathfrak{Z}\left(P\right)$. The first step of our algorithm is replacement of the leaves $\langle 1\rangle$ by their order series $\mathfrak{Z}\left(1\right)=\frac{x}{(1-x)^{2}}$. In the tree, going from top to bottom, the algebra homomorphism allows us to make the following replacements: a concatenation PQP*Q by $\mathfrak{Z}\left(P\right)*\mathfrak{Z}\left(Q\right)$ and $D(n)$ by $\mathfrak{Z}\left(D(n)\right)$ in $P$. Using Proposition 22 and Example 26, as well as linearity of $D$ and bilinearity of $\ast$, one evaluates the vertex operations until we reach the root to obtain
$D(D(D(D(1)*\mathfrak{Z}\left(1\right)*D(1)))*\mathfrak{Z}\left(3\right))$
$=D(D(D((\mathfrak{Z}\left(3\right)+2\mathfrak{Z}\left(4\right))*\mathfrak{Z}\left(1\right)*(\mathfrak{Z}\left(3\right)+2\mathfrak{Z}\left(4\right))))*\mathfrak{Z}\left(3\right))$
$=D(D(D(\mathfrak{Z}\left(7\right)+4\mathfrak{Z}\left(8\right)+4\mathfrak{Z}\left(9\right)))*\mathfrak{Z}\left(3\right))$
$=D(D(D(\mathfrak{Z}\left(7\right)))*\mathfrak{Z}\left(3\right))+4D(D(D(\mathfrak{Z}\left(8\right)))*\mathfrak{Z}\left(3\right))$
$+4D(D(D(\mathfrak{Z}\left(9\right)))*\mathfrak{Z}\left(3\right))$
This is an illustrative algorithm, remember that we are abusing notation: $\ast$ and $D$ for concatenation and ’itsari operations must be applied in their respective algebra. This is justified by the algebra homomorphism. Following the same procedure we continue
@@
6. Representability
The Wixárika posets were selected due to their topological properties. We now use the topological information to answer the inverse problem: Given a function $f(x)$, does there exist a poset $P$ whose order series coincides with $f(x)$? In this section, we present an algorithm to answer this question.
Definition 38.
@@ 6. Representability
Example 39.
Two posets are called Doppelgänger posets if they share the same order series, for example $\langle 1\rangle*(\langle 1\rangle\sqcup\langle 1\rangle\sqcup\langle 1\rangle)$ and $\langle 2\rangle\sqcup\langle 2\rangle$ have the same order series: $\mathfrak{Z}\left(2\right)+6\mathfrak{Z}\left(3\right)+6\mathfrak{Z}\left(4\right)$ [HPP+18].
@@
Theorem 40.
Let $P$ be a Wixárika poset. Consider its factorization, that is, the associated word $w$ in the characters * and $D$. Let $f(x)=w(\mathfrak{Z}\left(1\right))$ be the power series obtained by replacing every leaf of $w$ by $\mathfrak{Z}\left(1\right)$ and proceeding as in Example 37.
  1. (1)
@@
    where the coefficients $d_{j}$ are non-negative integers.
  2. (2)
@@
One shows that every Wixárika poset has a unique minimum and a unique maximum. Then all statements follow by describing the possible values obtained under the operations $D$ and *. The explicit description of $D$ and * leads to the positivity of the coefficients.
Consider, for example, statement $(4)$: We know that by applying $D$ we increase the first Betti number of the Hasse diagram of a poset by 1. The first Betti number of a concatenation is the sum of the first Betti numbers of each element, because Wixárika posets have a unique maximum and a unique minimum. Since Wixárika posets are generated by $\langle 1\rangle$, and the Betti number of $\langle 1\rangle$ is zero, it follows that the only way to increase the Betti number of a poset is by the operation $D$ or concatenating the poset with another poset with non-zero Betti number.
For details, see [ABD23, Proposition 3.1]. ∎
Example 41.
@@
The paper [BHK18] calls the elements $\{\mathfrak{Z}\left(i\right)\}_{i\in\mathbb{N}}$ the chain basis. The coefficients of order series in the chain basis satify the following positivity property:
Theorem 42.
@@
Sketch of the proof.
We first verify that the function $f(x)$ satisfies the requirements of Theorem 42. Then we restrict the search space using Theorem 40. Using colored operads, it is possible to reduce the search to a set $C$ of representatives of equivalence classes of trees up to associativity. Now for every representative $P\in C$ we first evaluate the first and last term of $\mathfrak{Z}\left(P\right)=b_{i}\mathfrak{Z}\left(i\right)+\cdots+b_{k}\mathfrak{Z}\left(k\right)$. And this evaluation can be optimized as well since we know what are the final values. If we find a poset $Q$ with $\mathfrak{Z}\left(Q\right)=r_{i}\mathfrak{Z}\left(i\right)+\cdots+r_{k}\mathfrak{Z}\left(k\right)$ and $b_{i}=r_{i},b_{k}=r_{k}$ then we compute all terms ${\{r_{j}\}}_{i<j<k}$ and compare with $f(x)$ until we find a solution of we finish the candidate posets. The evaluation can be done with the tree of operations as described in [FS86, Section 4], but working with power series instead of polynomials. If we find a poset that produces $f(x)$, we use the results of [BHK18, Sections 4.2 and 5 ] to find the remaining Doppelgängers, that is, posets that have the same power series $f(x)$. After this step, we stop the algorithm. For details, see [ABD23, Proposition 3.2]. ∎
@@
Remark 44.
Theorem 43 requires a function already written as a finite linear combination of terms $\{\mathfrak{Z}\left(n\right)\}_{n\in\mathbb{N}}$. The reason is that a random power series may not be expressable as a finite linear combination of the $\{\mathfrak{Z}\left(n\right)\}_{n\in\mathbb{N}}$. For example, consider the function $\sum_{n=1}^{\infty}\sum_{j=1}^{n}{n\choose j}x^{n}$.
Example 45.
Consider the function $8\mathfrak{Z}\left(13\right)+25\mathfrak{Z}\left(14\right)+18\mathfrak{Z}\left(15\right)$. The algorithm of Theorem 43 efficiently tests posets until it finds a candidate. Suppose the algorithm finds $Q=D(\mathfrak{Z}\left(8\right))*D(\mathfrak{Z}\left(1\right))$. Then with [BHK18, Section 5.3] we obtain $Q_{2}=D(D(\mathfrak{Z}\left(2\right)))*\mathfrak{Z}\left(7\right)$ as a Doppelgänger of $Q$. Up to commutativity of *, these are all the Doppelgänger Wixárika posets.
7. Farewell
@@
What other operads can interested readers study? There is an encyclopedia of operads, a project in progress by Frédéric Chapoton: Operadia https://operadia.pythonanywhere.com/table, and the Encyclopedia of types of algebras [ZIN10]. If one aims to study a new operad, the key point is to have a notion of composition, for example, cylinders can be composed vertically. A variation of the operad of cylinders encodes all the operations one expects in calculus, as described in [DTT08, Theorem 2].
The following list is not exhaustive, but a sample to illustrate papers with applications of operad theory. After all, in February 2026 the website zbMATH returns 264 papers with the word ”operad” in the title. Operad theory started with the work of J. M. Boardman, R. M. Vogt [BV68], and P. May [MAY72], in the field of algebraic topology. M. Kontsevich [KON99], D. Tamarkin [TAM03], and T. Willwacher [ŠW11] popularized them in mathematical physics. F., Chapoton [CHA07] described the relationship between operads and combinatorial species, S. Giraudo [GIR17] studies combinatorial problems as algebras of some operads. F. Fauvet, L. Foissy, and D. Manchon [FFM18] studied the operadic structures on finite posets. J. Baez, J. Foley, J. Moeller, and B. S. Pollard studied network models via operads in [BFM+20].
@@
8. Acknowledgments
An old version of these notes was used to give a minicourse in Govt. College University Lahore in Pakistan 2022, and at the BIRS-CMO workshop “6th meeting of the Mexican Mathematicians in the World” 2024. The work of E. Dolores Cuenca was supported by the Korea National Research Foundation (NRF) grant funded by the Korean government (MSIT) (RS-2025-00517727).

### v3

@@ 4 Algebras over the operad of Wixárika posets
A more detailed definition of operad can be found in  [23, 7, 17, 25, 19, 4, 14].
Example 4.3.
@@
Example 4.4.
Consider the tree grafting operad: Let $T(n)$ be the set of trees with $1$ root and $n$ leaves labeled (arbitrarily) $1$ through $n$. The collection $\{T(n)\}$ forms an operad by grafting the root of $g$ to the leaf of $f$ labeled $i$, as in Figure 4. Once the grafting is done, the leaves of $g$ are relabeled adding i1i-1 to the original label. Leaves of $f$ after the leaf $i$ are also relabeled continuing the numeration $i+m,\cdots,m+n-1$.
Figure 4: Operations on the operad tree. The tree $g$ is grafted in the tree $f$ using the leaf positioned at $i$ (Image created by Wiki_cies, https://commons.wikimedia.org/w/index.php?title=User:Wiki_cies&action=edit&redlink=1).
Definition 4.5.
@@
Definition 4.6.
Let $(P,\leq_{P})$ be a finite poset with cardinality $|P|=n$. Let $\mathbf{Q}=\{(Q_{i},\leq_{Q_{i}})\}_{i\in P}$ be a collection of finite posets indexed by the elements of $P$. The lexicographic sum of elements in the collection $\mathbf{Q}$ along $P$ is the poset with underlying set $\bigcup Q_{i}$ and the order $x\leq y$ is given by:
  1. 1.
    if $x,y\in Q_{i}$ and $x\leq_{Q_{i}}y$ or
  2. 2.
    if $x\in Q_{i}$, $y\in Q_{j}$, and $i\leq_{P}j$.
Example 4.7.
Recall that we defined the binary endomorphism $\ast$ on the set of finite posets in Definition 2.6. Given finite posets $Q_{x}$ and $Q_{y}$, the lexicographic sum of $Q_{x}$ and $Q_{y}$ along $\{x<y\}$ is $Q_{x}\ast Q_{y}$.
Definition 4.8.
The Wixárica operad is the operad $W$ generated under lexicographic sum by the unary operation $D$, the identity, and the binary operation $\ast$.
Example 4.9.
@@
Figure 5: Left: Trees illustrating the identity operation, the operation $D$ and the operation $\ast$. Right: A tree representing the element in Figure 3, which is obtained using only the operations $D$ and $\ast$.
The lexicographic sum then allows us to see Wixárika posets as algebras over the operad $W$ generated by points.
Example 4.10.
Consider Figure 6. On the left side we have at the top three copies of the posets $\langle 1\rangle$. Below is a sequence of transformations using $D$ and *. At the bottom we have the final result, a Wixárika poset. On the right side we have the corresponding tree of operations in which the leaves should be thought of as decorated by the poset $\langle 1\rangle$. The vertices with two incoming edges and one outgoing edge are labeled with * and the vertices with one incoming edge and one outgoing edge (the bird nest) are labeled with $D$. From top to bottom, as drops of water (the posets) located on the leaves fall due to gravity, the drops reach the branches labeled with operations. When passing through, drops are replaced by the output of the operations. This continues until we reach the root.
Figure 6: To the left bottom we have a Wixárika poset, as an element of the algebra. To the right we have the tree of operations, an element of the operad of Wixárika posets. When this tree of operations is evaluated on six points it produces the Wixárika poset of the left. As the drops of water run down the tree, the poset on the left is being assembled.
Definition 4.11.
@@
Proposition 4.12.
@@
This proposition is another way to paraphrase Equations (3.3) and (3.4). Put together Definition 3.7 and Proposition 3.10 to get:
$\mathfrak{Z}\left(P\ast Q\right)=\mathfrak{Z}\left(P\right)(1-x)\mathfrak{Z}\left(Q\right)=\mathfrak{Z}\left(P\right)\hat{*}\mathfrak{Z}\left(Q\right).$
By Proposition 3.11 and its Corollary 3.12 we have:
[…]
We obtain the explicit formula
[…]
5 Evaluation on order series
In this section, we give an explicit example of how to compute the order series of a Wixárica poset using the algebra homomorphims between both $W$-algebras defined in the previous section. To compute the order series, we follow [9], but work on the level of power series instead of order polynomials.
[…]
$\mathfrak{Z}\left(P\right)\hat{*}\mathfrak{Z}\left(Q\right)=\mathfrak{Z}\left(P*Q\right),\quad\widehat{D}(\mathfrak{Z}\left(P\right))=\mathfrak{Z}\left(D(P)\right).$
@@ 5 Evaluation on order series
We shall give a procedure to calculate $\mathfrak{Z}\left(P\right)$. The first step of our algorithm is replacement of the leaves $\langle 1\rangle$ by their order series $\mathfrak{Z}\left(1\right)=\frac{x}{(1-x)^{2}}$. In the tree, going from top to bottom, the algebra homomorphism allows us to make the following replacements: a concatenation PQP*Q by $\mathfrak{Z}\left(P\right)\hat{*}\mathfrak{Z}\left(Q\right)$ and $D$ by $\widehat{D}$. Using Proposition 3.10 and Example 3.14, as well as linearity of $\widehat{D}$ and bilinearity of $\hat{*}$, one evaluates the vertex operations until we reach the root to obtain
$\widehat{D}\left(\widehat{D}\left(\widehat{D}\left(\widehat{D}(1)\hat{*}\mathfrak{Z}\left(1\right)\hat{*}\widehat{D}(1)\right)\right)\hat{*}\mathfrak{Z}\left(3\right)\right)$
$=\widehat{D}\left(\widehat{D}\left(\widehat{D}\left(\left(\mathfrak{Z}\left(3\right)+2\mathfrak{Z}\left(4\right)\right)\hat{*}\mathfrak{Z}\left(1\right)\hat{*}\left(\mathfrak{Z}\left(3\right)+2\mathfrak{Z}\left(4\right)\right)\right)\right)\hat{*}\mathfrak{Z}\left(3\right)\right)$
$=\widehat{D}\left(\widehat{D}\left(\widehat{D}\left(\mathfrak{Z}\left(7\right)+4\mathfrak{Z}\left(8\right)+4\mathfrak{Z}\left(9\right)\right)\right)\hat{*}\mathfrak{Z}\left(3\right)\right)$
$=\widehat{D}\left(\widehat{D}\left(\widehat{D}(7)\right)\hat{*}\mathfrak{Z}\left(3\right)\right)+4\widehat{D}\left(\widehat{D}\left(\widehat{D}(8)\right)\hat{*}\mathfrak{Z}\left(3\right)\right)+4\widehat{D}\left(\widehat{D}\left(\widehat{D}(9)\right)\hat{*}\mathfrak{Z}\left(3\right)\right).$
Following the same procedure we continue:
@@
6 Representability
We now use the topological information to answer the inverse problem: Given a function $f(x)$, does there exist a Wixárika poset $P$ whose order series coincides with $f(x)$? In this section, we present an algorithm to answer this question.
Definition 6.1.
@@ 6 Representability
Example 6.2.
Two posets are called Doppelgänger posets if they share the same order series, for example $\langle 1\rangle*(\langle 1\rangle\sqcup\langle 1\rangle\sqcup\langle 1\rangle)$ and $\langle 2\rangle\sqcup\langle 2\rangle$ have the same order series: $\mathfrak{Z}\left(2\right)+6\mathfrak{Z}\left(3\right)+6\mathfrak{Z}\left(4\right)$ [13].
@@
Theorem 6.3.
Let $P$ be a Wixárika poset. Consider its factorization, that is, the associated word $w$ in the characters * and $D$. Let $f(x)=w(\mathfrak{Z}\left(1\right))$ be the power series obtained by replacing every leaf of $w$ by $\mathfrak{Z}\left(1\right)$ and proceeding as in Example 5.1.
  1. 1.
@@
    where the coefficients $d_{j}$ are positive integers.
  2. 2.
@@
All statements follow by describing the possible values obtained under the operations $\widehat{D}$ and $\hat{*}$. The explicit description of $\widehat{D}$ and $\hat{*}$ leads to the positivity of the coefficients.
Consider, for example, statement $(4)$: Our goal is to show that the difference kik-i counts the first Betti number of the Hasse diagram of the poset. We can argue inductively:
The first Betti number of $\langle 1\rangle$ is zero, and ki=0k-i=0 for $\mathfrak{Z}\left(1\right)$.
We know that by applying $D$ we increase the first Betti number of the Hasse diagram of a poset by 1. We also know the precise effect of applying $\widehat{D}$ on $\mathfrak{Z}\left(n\right)$. According to Example 3.14 the difference kik-i is increased by 1. By linearity (Proposition 3.13) the effect of $\widehat{D}$ on the order series of a poset increases kik-i by 1 as well.
Wixárika posets satisfy the following property: The first Betti number of the poset obtained by concatenation is the sum of the first Betti numbers of the input posets. This additivity also holds for the difference kik-i: Given two power series of Wixárika posets $f,g$, from Proposition 3.7 we see that $f\hat{*}g$ satisfies $(k_{f}+k_{g})-(i_{f}+i_{g})=(k_{f}-i_{f})+(k_{g}-i_{g})$.
Finally, since Wixárika posets are generated by $\langle 1\rangle$ under the operations D,D,*, we conclude that the difference kik-i counts the first Betti number of a Wixárika poset.
For details, see [1, Proposition 3.1]. ∎
Example 6.4.
@@
The paper [6] calls the elements $\{\mathfrak{Z}\left(i\right)\}_{i\in\mathbb{N}}$ the chain basis. The coefficients of order series in the chain basis satify the following positivity property:
Theorem 6.5.
@@
Sketch of proof.
Given $f$ we test posets to see whether $\mathfrak{Z}\left(P\right)=f$ for some poset $P$. We use geometric information to reduce the search space.
First, we verify that $f$ satisfies conditions (1) and (7) of Theorem 6.3. Using the remaining conditions of Theorem 6.3, we then filter candidate posets (e.g., Betti number equal to kik-i, number of points equal to $i$, etc). Next, we pass to equivalence classes of posets up to associativity (for example, by using colored operads to factorize the posets). We then evaluate candidate posets in stages. For every representative $P$ of an equivalence class of posets, we only evaluate the first and last element in the expansion $\mathfrak{Z}\left(P\right)=\sum_{j=i}^{k}c_{j}\mathfrak{Z}\left(j\right)$, that is, we keep track only of $c_{i}\mathfrak{Z}\left(i\right)$ and $c_{k}\mathfrak{Z}\left(k\right)$.
At every stage of the evaluation of $\mathfrak{Z}\left(P\right)$, if we obtain partial coefficients $(c_{i}^{\prime}\mathfrak{Z}\left(i\right),c_{k}^{\prime}\mathfrak{Z}\left(k\right))$, we require that $c_{i}^{\prime}|d_{i}$, otherwise the poset can be discarded. The evaluation is performed using the tree of operations as described in [9, Section 4], but working with power series instead of polynomials.
If we find a poset $Q$ with $\mathfrak{Z}\left(Q\right)=r_{i}\mathfrak{Z}\left(i\right)+\cdots+r_{k}\mathfrak{Z}\left(k\right)$ and $r_{i}=d_{i},r_{k}=d_{k}$ then we compute all intermediate coefficients ${\{r_{j}\}}_{i<j<k}$ and compare the result with $f(x)$. We continue until we find a solution or all candidate posets have been tested. If a poset producing $f(x)$ is found, we use the results of [6, Sections 4.2 and 5 ] to determine all Doppelgängers, that is, posets with the same power series $f(x)$. After this step, the algorithm terminates. For details, see [1, Proposition 3.2]. ∎
@@
Remark 6.7.
The results of [6, Sections 4.2 and 5 ] rely on the values of the coordinates in the binomial basis, while Proposition 6.3 rely on the position of the non zero coordinates in the binomial basis. In that sense, both results are complementary to each other.
Remark 6.8.
Theorem 6.6 requires a function already written as a finite linear combination of terms $\{\mathfrak{Z}\left(n\right)\}_{n\in\mathbb{N}}$. The reason is that a random power series may not be expressable as a finite linear combination of the $\{\mathfrak{Z}\left(n\right)\}_{n\in\mathbb{N}}$. For example, consider the function $\sum_{n=1}^{\infty}\sum_{j=1}^{n}{n\choose j}x^{n}$.
Example 6.9.
Consider the function $8\mathfrak{Z}\left(13\right)+25\mathfrak{Z}\left(14\right)+18\mathfrak{Z}\left(15\right)$. The algorithm of Theorem 6.6 efficiently tests posets until it finds a candidate. Suppose the algorithm finds Q=D(8)D(1)Q=D(8)*D(1). Then with [6, Section 5.3] we obtain $Q_{2}=D(D(2))*\langle 7\rangle$ as a Doppelgänger of $Q$. Up to commutativity of *, these are all the Doppelgänger Wixárika posets.
7 Farewell
@@
What other operads can interested readers study? There is an encyclopedia of operads, a project in progress by Frédéric Chapoton: Operadia https://operadia.pythonanywhere.com/table, and the Encyclopedia of types of algebras [28] by Jean-Louis Loday. If one aims to study a new operad, the key point is to have a notion of composition, for example, cylinders can be composed vertically. A variant of the operad of cylinders encodes all the operations one expects in calculus, as described in [8, Theorem 2].
The following list is not exhaustive, but a sample to illustrate papers with applications of operad theory. After all, in February 2026 the website zbMATH returns 264 papers with the word “operad” in the title. Operad theory started with the work of J. M. Boardman, R. M. Vogt [3], and P. May [18] in the field of algebraic topology. M. Kontsevich [16], D. Tamarkin [24], and T. Willwacher [21] popularized them in mathematical physics. F. Chapoton [7] described the relationship between operads and combinatorial species, S. Giraudo [11] studied combinatorial problems as algebras over operads. F. Fauvet, L. Foissy, and D. Manchon [10] studied the operadic structures on finite posets. J. Baez, J. Foley, J. Moeller, and B. S. Pollard studied network models via operads  [2].
@@
Acknowledgments
An old version of this note was used to give a minicourse in Govt. College University Lahore in Pakistan 2022, and at the BIRS-CMO workshop “6th meeting of the Mexican Mathematicians in the World” 2024. The work of E. Dolores Cuenca was supported by the Korea National Research Foundation (NRF) grant funded by the Korean government (MSIT) (RS-2025-00517727).